Alex drove 270 miles to the University of Phoenix (UOP) in the same time as Stephanie drove 330 miles to UOP. If Stephanie drove 10 miles per hour faster than Alex, then how fast did each of them drive?
Answer by mananth(16949)

Alex speed = xmph: distance = 270 miles
Stephanie speed = x+10 mph: 330 miles
..
time Alex= time Stephanie
t=d/r
270/x=330/(x+10)
cross multiply
270(x+10)=330x
270x+2700=330x
330x-270x=2700
30x=2700
/30
x=90 mph : speed Alex
90+10 = 100 ... speed ... Stephanie
...
m.ananth@hotmail.ca

..